# Enhancing Facial Expression Recognition through Dual-Direction Attention Mixed Feature Networks: Application to 7th ABAW Challenge

### v2

@@ 3 Results
Overall, both approaches—fine-tuning the entire DDAMFN model and training a custom classifier—performed well across the evaluated metrics, with fine-tuning the DDAMFN model showing slightly better performance in individual task metrics.
@@ 4 Conclusion
The results using DDAMFN as a feature extractor with the pretrained weiths are practically the same as those obtained by fine-tuning the entire model. This similarity in performance can be attributed to the architecture of DDAMFN, which has been trained to generalize exceptionally well for the multitask challenge problem. This inherent capability allows it to extract meaningful features that are sufficient for achieving high performance without the need for extensive fine-tuning.
In contrast, the single-task experiment revealed a different scenario. Fine-tuning the entire model for the single-task yielded better results than merely training the classifier, and even outperformed the multitask results. This finding underscores the importance of task-specific optimization. When the model is fine-tuned for a specific task, it can leverage the nuances and particularities of that task, leading to superior performance.

### v3

@@ 3 Results
Overall, both approaches: fine-tuning the entire DDAMFN model and training a custom classifier, performed well across the evaluated metrics, with fine-tuning the DDAMFN model showing slightly better performance in individual task metrics.
4 Test set evaluation
We conducted an extensive evaluation by submitting three distinct sets of prediction models. The first set of predictions was generated using the model that had been fine-tuned across the entire architecture (DDAMFN+thresholds).
The second set of predictions was done leveraging DDAMFN as a feature extractor. In this approach, the sophisticated feature extraction capabilities of DDAMFN were utilized to capture complex patterns and representations within the data. A custom classifier was then trained on these extracted features, allowing for a focused and potentially more accurate classification performance(Custom classifier+thresholds).
The third set of predictions was derived from the ensemble approach, where the results from the models trained individually for each task were combined.
| Models for each submission | P | |||
| DDAMFN+thresholds | 0.262 | 0.282 | 0.478 | 1.022 |
| Custom classifier+thresholds | 0.371 | 0.277 | 0.466 | 1.114 |
| Ensambled individual model results | 0.292 | 0.272 | 0.499 | 1.064 |
Results are detailed in Table 5. DDAMFN+thresholds: This model achieved the highest F1 score for expressions ( = 0.282), indicating its effectiveness in classifying expressions. However, it had the lowest combined CCC for valence and arousal ( = 0.262), suggesting it is less reliable in predicting emotional valence and arousal. The overall performance metric P was 1.022, positioning it as the least effective in terms of the combined performance metrics.
Custom classifier+thresholds: This model excelled with the highest combined CCC for valence and arousal ( = 0.371), showing its strength in predicting emotional states accurately. It also had the highest overall performance metric P (1.114), reflecting its robust performance across the evaluated metrics. Despite this, its F1 scores for expressions ( = 0.277) and action units ( = 0.466) were slightly lower compared to the other models, indicating room for improvement in those specific areas.
Ensembled individual model results: This model achieved the highest F1 score for action units ( = 0.499), making it the best performer in classifying action units. Its combined CCC for valence and arousal was moderate ( = 0.292), and its F1 score for expressions was the lowest among the models ( = 0.272). The performance metric P was 1.064, indicating a balanced performance but not leading in any single category except for action units.
The model that achieved the highest overall performance was ranked 5th in the 7th Affective Behavior Analysis in the Wild (ABAW) Challenge. This ranking reflects its competitive capability and robustness among numerous submissions from diverse research groups, highlighting its efficacy in real-world affective behavior analysis tasks.
| Teams | P | |||
| Netease Fuxi AI Lab [23] | 0.5420 | 0.4286 | 0.5580 | 1.5286 |
| HSEmotion [29] | 0.4074 | 0.3279 | 0.5119 | 1.2472 |
| HFUT-MAC1 [30] | 0.3783 | 0.2997 | 0.4997 | 1.1777 |
| SCU-ACers [22] | 0.3743 | 0.3018 | 0.4879 | 1.1640 |
| Ours | 0.3710 | 0.2772 | 0.4643 | 1.1145 |
| SML | 0.292 | 0.1938 | 0.4046 | 0.8692 |
| baseline | 0.1193 | 0.1007 | 0.1200 | 0.3400 |
The key differences compared to other teams are multifaceted. Primarily, other teams have employed more sophisticated and heavier feature extractors, such as Masked Autoencoders (MAE) [23], or have integrated multiple feature extractors to enhance the overall quality of the extracted features [30]. This approach has allowed them to achieve superior results.
In contrast, our methodology included a more aggressive data filtering process, which consequently led to the utilization of a smaller dataset for training purposes compared to our competitors. This reduction in data volume may have impacted our model’s ability to generalize effectively.
Moreover, a significant limitation of our training approach has been the treatment of each frame in isolation, without taking into account any temporal relationships or dependencies within the sequence of frames [29]. This lack of temporal context has posed challenges for the training process, potentially hindering the model’s performance in tasks where temporal dynamics are crucial.
Another notable distinction lies in our reliance on a single, streamlined model architecture. Unlike some other teams, which have explored the benefits of ensemble methods or hybrid models incorporating both classical and deep learning techniques, our approach has remained relatively straightforward [22].
Overall, while our methods have certain strengths, these differences in feature extraction, data volume, and temporal context consideration have been critical factors influencing our comparative performance.
@@ 5 Conclusion
The results using DDAMFN as a feature extractor with the pretrained weigths were practically the same as those obtained by fine-tuning the entire model. This similarity in performance can be attributed to the architecture of DDAMFN, which has been trained to generalize exceptionally well for the multitask challenge problem. This inherent capability allows it to extract meaningful features that are sufficient for achieving high performance without the need for extensive fine-tuning.
In contrast, the single-task experiment revealed a different scenario. Fine-tuning the entire model for the single-task yielded better results than merely training the classifier, and even outperformed the multitask results. This finding underscored the importance of task-specific optimization. When the model is fine-tuned for a specific task, it can leverage the nuances and particularities of that task, leading to superior performance.
@@
Interestingly, although the least performant model on the evaluation set was the Custom classifier + thresholds, it achieved the best score on the test set. This suggests that the Custom classifier + thresholds might have better generalization capabilities, highlighting the importance of thorough evaluation on diverse test sets.
@@
Acknowledgments
This work is part of the project PLEC2021-007868, funded by MCIN and by the EU, and the project PID2022-138721NB-I00, funded by MCIN.